Calculate the wave length s of the following cases : 1)neutrons thermally excited at T = 500 degrees K 2) electrons accelerated by a voltage of 50kV.
1
Expert's answer
2020-07-20T06:21:21-0400

Answer.

First part

Energy of neutron (E)= 3kT/2 = 1/2mv2

V = (3kT/m)1/2

Wavelength of the neutron = h/mv = h/(3mkT)1/2

m = mass of neutron = 1.67 x 10-27 kg

h = Plank constant = 6.627 x 10-34 JS

k = Boltzman constant = 1.38 x 10-23 JK-1

Putting these values, wave length of neutron at 500K is = 1.127 x 10-10 m (Ans)

Second part

Charge of an electron = 1.6 x 10-19 C

Mass of the electron (m) = 9.11 x 10-31 kg

Kinetic energy of the electron (E)= 1.6 x 10-19  x 50 x 1000 J = 8 x 10-15 J

Wavelength of the electron = h/(2mE)1/2

h = Plank constant = 6.627 x 10-34 JS

Putting these values, wavelength of the electron = 5.467 x 10-12 m (Ans.)